# Reinforced Generation of Combinatorial Structures: Hardness of Approximation

### v5

@@ 5 Discussion on AI-assisted Mathematics and Complexity Theory
The best test of robustness for AI-assisted mathematical discovery is durability: whether or not new results such as ours are superseded by humans (possibly with computer assistance). We note that some results in AI-assisted mathematics have in fact been matched or improved quickly without the use of AI [19, 9], and in some cases the results existed (even before the problem was posed [47]) but were not known to the authors when AI methods were applied [2].
@@
Third, our work suggests that gadget-based reductions lend themselves to optimization beyond traditional methods (e.g., MIP/SMT solvers), using AlphaEvolve. This in turn suggests that beating AlphaEvolve will generally require non-gadget methods like custom PCPs [3].
Finally, it is worth dwelling on some failures of our approach. For some problems even if verification is trivial we could not get AlphaEvolve to work. An example of this is the Hadamard-$668$ conjecture which states that there exists a Hadamard matrix of dimensions $668\times 668$. (More generally, it is conjectured that a Hadamard matrix $H_{n\times n}$ exists for any $n$ which is a multiple of four; 668 is the smallest value for which this is not known.) We attempted to construct one using AlphaEvolve. Although the search is over a large space of $2^{668^{2}}$ possibilities, verifying any candidate for correctness only takes $2,23,112$ bitwise multiplications. Even with fast verification, AlphaEvolve was unable to find a construction for $H_{668\times 668}$. In fact, AlphaEvolve even failed to replicate the construction for $H_{428\times 428}$, which was previously the smallest order for which no construction was known until [34]; this despite the fact that the construction for 428 is publicly available on the internet.
Looking ahead, it is conceivable that advances in LLM reasoning [38, 24, 51] could be coupled with AlphaEvolve, especially in generating the initial code-snippet and more effective problem specific prompting of the LLMs used by AlphaEvolve. It is possible that our methods can be extended to other problems in extremal combinatorics. We leave exploration of these problems as future directions.
@@ 6 Acknowledgments
We thank Adam Zsolt Wagner for helping us with AlphaEvolve throughout this project, and for his invaluable advice on our experimental setup — his insights speeded up our experimentation significantly. Swarat Chaudhuri not only helped us with our initial work on MAX-3-CUT, but also generously advised us throughout on various approaches to verification — we are deeply grateful to him for this. Sushant Sachdeva worked closely with us on MAX-3-CUT, TSP, and the design of verifiers and freely shared his valuable intuition on gadget reductions. Pasin Manurangsi replicated some of our early gadgets by hand (especially the optimal gadget that reduces from $P_{0}^{=}$ to $P^{\neq}$), boosting our confidence early on that AlphaEvolve was progressing in the right direction. Pasin also helped us solidify our understanding of [33, 11]. Sidhanth Mohanty made critical contributions in improving the upper bounds on average-case hardness (Theorem 2.2). Krishnamurthy (Dj) Dvijotham helped us with the advanced use of MIP solvers, and ablation studies with some of our verifiers. His ablation studies convinced us that direct usage of MIP solvers would not scale to the gadget sizes we are operating with. Shuang Song helped us with numerous engineering challenges. We thank Raluca Ada Popa, Mary Chesus, Jonathan Katz, Ravi Kumar, James Manyika, Yossi Matias, Jelani Nelson, Rina Panigrahy, Amit Sahai, Thomas Steinke and Jalaj Upadhyay for their valuable feedback on the manuscript. We thank Uri Feige for pointing out an error in a previous version of Appendix B.2. We thank Four Flynn and Pushmeet Kohli for their continued support and feedback through the course of this project.
@@ Appendix A AlphaEvolve as a Framework for Combinatorial Discovery
A.1 Propose-test-refine (PTR) paradigm
We operate in the PTR paradigm  [^10] The name is motivated from the Propose-test-release (PTR) framework in the differential privacy literature [15], which performs a similar task of testing a proposed candidate construction for (privacy) properties, before releasing it. for setting up experiments to discover novel combinatorial structures. While the paradigm is implicit in prior work [42], we first make it explicit, and then instantiate the extensions needed in this paper. The PTR paradigm consists of three components:
@@ A.1 Propose-test-refine (PTR) paradigm
    AlphaEvolve makes an LLM call using the tuple of inputs in the prompt to the LLM, prompting it to output the description of $\mathcal{C}_{t+1}$, with the goal of increasing the score [^12] Technically, AlphaEvolve does not feed in all the prior functions as the prompt to the LLM. Instead, it uses a genetic algorithm to select a subset of them while maintaining high scores, and diversity [48, Figure 1].. We emphasize the following subtle point: AlphaEvolve does not directly search for combinatorial structures, but instead for a code snippet that would output a combinatorial structure with a high score. Figure 6 shows an initial code-snippet to be evolved by AlphaEvolve.
@@
    Our contributions: Our main contributions to the PTR paradigm are the following: a) using AlphaEvolve itself to accelerate the verification for $\textup{MAX-}{k}\textup{-CUT}$ (sometimes up to $10,000\times$), and b) an approach towards using AlphaEvolve for solving optimization problems that involve mixed-integer programming (MIP) style constraints. For the $k=3$ case, the AlphaEvolve optimized verification code constructed a solver that offloaded a $O(3^{m})$ time computation to a highly optimized tensor contraction operation in numpy, and only performing slow python-based computation for $O(3^{m/3})$ time. We discussed each of these contributions in detail in Sections 3.1 and 3.2.
@@ Appendix B Proofs of Average Case Theorems
For all $d\in\mathbb{N}$ $\varepsilon>0$, with probability $1-o_{n}(1)$ over $G\sim\mathcal{G}(n,d)$, we have $\lambda^{*}(G)\leq 2\sqrt{d-1}+\varepsilon$.
@@
The $d$-ary tree of depth $L$ $T_{d,L}$ is a rooted tree, where the root has $d$ children, and all non-root vertices at distance <L<L from the root have d1d-1 children.
@@
We will deduce some linear constraints on $\mu_{G,S}$, in terms of its pdf.
@@
An interesting point is that the $d=4,L=3$ case has $11880$ variables and is technically within reach of current solvers, but resulted in the same bound on $\sigma^{\textup{MC}}_{4}$ as the $d=4,L=2$ case.
@@
In this section, we give the details of the NP-hardness of approximation proofs for MAX-k-CUT for $k\in\{3,4\}$. For ease of exposition, we begin by reintroducing all predicates needed for the proof. All predicates will be over the alphabet $\Psi=\mathbb{Z}_{k}$.
  • $P^{\neq}$ denotes a $2$-ary predicate defined by $P^{\neq}(x,y)=\mathbf{1}\{x\neq y\}$.
  • For $i\in\{0,\ldots,k-1\}$, $P^{=}_{i}$ denotes a $3$-ary predicate defined by $P^{=}_{i}(x,y,z)=\mathbf{1}\{x+y+z\equiv i\pmod{k}\}$.
[…]
@@
Let $k\geq 2$ and define $\Psi=\mathbb{Z}_{k}$.
  • $i$-gadget: For $i\in\Psi$, an $i$-gadget from $3\textup{LIN}({k})$ to $\textup{MAX-}{k}\textup{-CUT}$ is an instance $I_{i}$ of $\textup{MAX-}{k}\textup{-CUT}$ over $3+k+n_{\textup{aux}}$ variables, where $n_{\textup{aux}}\in\mathbb{N}$.
  • Primary, global, and auxiliary variables: We partition the set of variables $3+k+n_{\textup{aux}}]$ into three primary variables $\{1,2,3\}$, $k$ global variables $\{4,5,\ldots,k+3\}$, and $n_{\textup{aux}}$ auxiliary variables $\{k+4,\ldots,3+k+n_{\textup{aux}}\}$. We will always write an assignment to the variables as a tuple $(\mathbf{x},\mathbf{y},\mathbf{z})\in\Psi^{3+k+n_{\textup{aux}}}$ where $\mathbf{x}\in\Psi^{3}$, $\mathbf{y}\in\Psi^{k}$, and $\mathbf{z}\in\Psi^{n_{\textup{aux}}}$.
The variables $\mathbf{y}$ will be restricted in a way to break the symmetries of $\textup{MAX-}{k}\textup{-CUT}$. Concretely, let $Y$ be the set of strings $\mathbf{y}$ in $\Psi^{k}$ that are lexicographically not larger than any string $\mathbf{y}^{\sigma}=(\sigma(y_{1}),\sigma(y_{2}),\ldots,\sigma(y_{k}))$ that is obtained by permuting the alphabet $\Psi$ by some permutation $\sigma:\Psi\to\Psi$. In particular, note that the vector $(0,1,2,\ldots,k-1)$ is in $Y$. We will only ever assign $\mathbf{y}$ to values in $Y$.
@@
We will set $C_{i}=(P_{i}^{=})^{-1}(1)$ and $S_{i}=(P_{i}^{=})^{-1}(0)$ to be the set of satisfying and unsatisfying assignments for the source predicate $P_{i}^{=}$ respectively.
We will associate four real-valued parameters to the gadget $I_{i}=\sum_{j\in[m]}\zeta_{j}$ as follows.
  • (Completeness analysis) $c(I_{i})$ is defined as
    $c(I_{i})=\min_{\mathbf{x}\in C_{i}}\max_{\mathbf{z}\in\Psi^{n_{\textup{aux}}}}I_{i}(\mathbf{x},\mathbf{y},\mathbf{z})$
    for $\mathbf{y}=(0,1,\ldots,k-1)$. For $\mathbf{x}\in C_{i}$, we say that $\arg\max_{\mathbf{z}\in\Psi^{n_{\textup{aux}}}}I_{i}(\mathbf{x},\mathbf{y},\mathbf{z})$ is the witness for $\mathbf{x}$.
  • (Soundness analysis) $s_{1}(I_{i})$ is defined as
    $s_{1}(I_{i})=\max_{\mathbf{x}\in C_{i},\mathbf{z}\in\Psi^{n_{\textup{aux}}},\mathbf{y}\in Y}I_{i}(\mathbf{x},\mathbf{y},\mathbf{z}).$
    Also, $s_{2}(I_{i})$ is defined as
    $s_{2}(I_{i})=\max_{\mathbf{x}\in S_{i},\mathbf{z}\in\Psi^{n_{\textup{aux}}},\mathbf{y}\in Y}I_{i}(\mathbf{x},\mathbf{y},\mathbf{z}).$
  • (Number of clauses) $t(I_{i})$ is defined to be $m$, that is, the number of clauses present in $I_{i}$.
@@
Let $k\geq 2$. Given an instance $I=\sum_{j\in[m]}\zeta_{j}$ of $3\textup{LIN}({k})$ with $m/k$ clauses from each of the predicates $P_{0}^{=},P_{1}^{=},\ldots,P_{k-1}^{=}$, it is NP-hard to distinguish between the following cases for any $\varepsilon>0$:
@@
    Completeness case. $\max_{\mathbf{x}}I(\mathbf{x})\geq(1-\varepsilon)\cdot m$, or
@@
    Soundness case. $\max_{\mathbf{x}}I(\mathbf{x})\leq(1/k+\varepsilon)\cdot m$.
@@
Let $k\geq 2$, and set $\Psi=\mathbb{Z}_{k}$. Let $I_{i}$ be an $i$-gadget from $3\textup{LIN}({k})$ to $\textup{MAX-}{k}\textup{-CUT}$ for each $i\in\Psi$. For any $\varepsilon>0$, given an instance $J$ of $\textup{MAX-}{k}\textup{-CUT}$ with $M$ constraints, it is NP-hard to distinguish between the following two cases for any $\varepsilon>0$:
@@
    Completeness case: $\max_{\mathbf{x}}J(\mathbf{x})\geq(a-\varepsilon)\cdot M$, or
@@
    Soundness case: $\max_{\mathbf{x}}J(\mathbf{x})\leq(b+\varepsilon)\cdot M$,
@@
$a=\frac{\sum_{i\in\Psi}c(I_{i})}{\sum_{i\in\Psi}t(I_{i})},\quad b=\frac{s_{1}(I_{r_{0}})+\sum_{i\in\Psi\setminus\{0\}}s_{2}(I_{r_{i}})}{\sum_{i\in\Psi}t(I_{i})}.$
Here $r_{0},r_{1},\ldots r_{k-1}$ is a permutation of $\{0,1,\ldots,k-1\}$ such that $s_{1}(I_{r_{i}})-s_{2}(I_{r_{i}})$ is maximized by $r_{0}$.
@@
We will perform a reduction starting from the NP-hard problem in Theorem C.1 to $\textup{MAX-}{k}\textup{-CUT}$. Let $I=\sum_{j\in[m]}\zeta_{j}$ be an $n$-variable instance of $3\textup{LIN}({k})$, and suppose that it has $m/k$ clauses using $P_{i}^{=}$ for all $i\in\Psi$. We will show how to reduce $I$ to an instance $J$ of $\textup{MAX-}{k}\textup{-CUT}$.
Let $n_{\textup{aux}}$ be the number of auxiliary variables in all gadgets $I_{i}$; we can assume they all have the same number of auxiliary variables by introducing new dummy variables to some of them. $J$ will be defined as an instance on $n+k+n_{\textup{aux}}\cdot m$ variables, which we will denote as $\mathbf{x}\in\Psi^{n}$, $\mathbf{y}\in\Psi^{k}$, and $\mathbf{z}\in\Psi^{n_{\textup{aux}}\cdot m}$. For $j\in[m]$ we will denote the $j^{th}$ block of size $n_{\textup{aux}}$ in $z$ as $z^{j}$. We will refer to $\mathbf{x},\mathbf{y},\mathbf{z}$ as the primary, global, and auxiliary variables respectively.
Let $\zeta_{j}=P^{=}_{i}(x_{j_{1}},x_{j_{2}},x_{j_{3}})$ be a clause corresponding to the predicate $P^{=}_{i}$ for some $i$. We will add a copy of $I_{i}$ on the variables $(\mathbf{x}^{\prime},\mathbf{y},\mathbf{z}^{j})$, where $\mathbf{x}^{\prime}=(x_{j_{1}},x_{j_{2}},x_{j_{3}})$. Note that the number of clauses in $J$ is $M=(m/k)\cdot\sum_{i\in\Psi}t(I_{i})$.
@@
    Completeness case: Assuming there is an $\mathbf{x}\in\Psi^{n}$ such that $I(\mathbf{x})\geq(1-\varepsilon)\cdot m$, we show how to construct an assignment $(\mathbf{x},\mathbf{y},\mathbf{z})\in\Psi^{n+k+n_{\textup{aux}}\cdot m}$ such that $J(\mathbf{x},\mathbf{y},\mathbf{z})\geq a\cdot M$. We will set $\mathbf{y}=(0,1,2,\ldots,k-1)$, and for each clause $\zeta_{j}=P_{i}^{=}(x_{j_{1}},x_{j_{2}},x_{j_{3}})$, $\mathbf{z}^{j}$ will be set as the witness $\mathbf{z}^{j}=\arg\max_{\mathbf{z}\in\Psi^{n_{\textup{aux}}}}I_{i}(\mathbf{x}^{\prime},\mathbf{y},\mathbf{z})$, where $\mathbf{x}^{\prime}=(x_{j_{1}},x_{j_{2}},x_{j_{3}})$.
@@
    $\sum_{i\in\Psi}c(I_{i})\cdot m/k-O(\varepsilon\cdot m)=a\cdot M-O(\varepsilon)\cdot m\geq(a-\varepsilon^{\prime})\cdot M$

### v6

@@ 5 Discussion on AI-assisted Mathematics and Complexity Theory
In the spirit of Turing’s imitation game [60], one strong test of robustness for AI-assisted mathematical discovery is durability: whether or not new results obtained with AI assistance are superseded by humans (possibly with computer assistance). We note that some results in AI-assisted mathematics have in fact been matched or improved quickly without the use of AI [23, 13], and in some cases the results existed (even before the problem was posed [3, 53]) but were not known to the authors when AI methods were applied [4, 9].
@@
Third, our work suggests that gadget-based reductions lend themselves to optimization beyond traditional methods (e.g., SMT/MIP solvers), using AlphaEvolve. This in turn suggests that beating AlphaEvolve will generally require non-gadget methods like custom PCPs [5].
Finally, it is worth dwelling on some failures of our approach. For some problems even if verification is trivial we could not get AlphaEvolve to work. An example of this is the Hadamard-$668$ conjecture which states that there exists a Hadamard matrix of dimensions $668\times 668$. (More generally, it is conjectured that a Hadamard matrix $H_{n\times n}$ exists for any $n$ which is a multiple of four; 668 is the smallest value for which this is not known.) We attempted to construct one using AlphaEvolve. Although the search is over a large space of $2^{668^{2}}$ possibilities, verifying any candidate for correctness only takes $2,23,112$ bitwise multiplications. Even with fast verification, AlphaEvolve was unable to find a construction for $H_{668\times 668}$. In fact, we failed to get AlphaEvolve to replicate the construction for $H_{428\times 428}$, which was previously the smallest order for which no construction was known until [40]; this despite the fact that the construction for 428 is publicly available on the internet.
Looking ahead, it is conceivable that advances in LLM reasoning [44, 29, 62] could be coupled with AlphaEvolve, especially in generating the initial code-snippet and more effective problem specific prompting of the LLMs used by AlphaEvolve. We leave exploration of these problems as future directions.
@@ 6 Acknowledgments
We thank Adam Zsolt Wagner for helping us with AlphaEvolve throughout this project, and for his invaluable advice on our experimental setup — his insights speeded up our experimentation significantly. Swarat Chaudhuri not only helped us with our initial work on MAX-3-CUT, but also generously advised us throughout on various approaches to verification — we are deeply grateful to him for this. Sushant Sachdeva worked closely with us on MAX-3-CUT, TSP, and the design of verifiers and freely shared his valuable intuition on gadget reductions. Pasin Manurangsi replicated some of our early gadgets by hand (especially the optimal gadget that reduces from $\texttt{P}^{\equiv 0}_{3,3}$ to $\texttt{P}^{\neq}_{2}$), boosting our confidence early on that AlphaEvolve was progressing in the right direction. Pasin also helped us solidify our understanding of [39, 15]. Sidhanth Mohanty made critical contributions in improving the upper bounds on average-case hardness (Theorem 2.2). Krishnamurthy (Dj) Dvijotham helped us with the advanced use of MIP solvers, and ablation studies with some of our verifiers. His ablation studies convinced us that direct usage of MIP solvers would not scale to the gadget sizes we are operating with. Shuang Song helped us with numerous engineering challenges. We thank Mary Chesus, Jonathan Katz, Ravi Kumar, James Manyika, Yossi Matias, Jelani Nelson, Rina Panigrahy, Raluca-Ada Popa, Amit Sahai, Thomas Steinke and Jalaj Upadhyay for their valuable feedback on the manuscript. We thank Uri Feige for pointing out an error in a previous version of Appendix B.2. We thank Four Flynn and Pushmeet Kohli for their continued support and feedback through the course of this project.
@@ Appendix A AlphaEvolve as a Framework for Combinatorial Discovery
Propose-test-refine (PTR) paradigm: We operate in the PTR paradigm  [^10] The name is motivated from the Propose-test-release (PTR) framework in the differential privacy literature [19], which performs a similar task of testing a proposed candidate construction for (privacy) properties, before releasing it. for setting up experiments to discover novel combinatorial structures. While the paradigm is implicit in prior work [48], we first make it explicit, and then instantiate the extensions needed in this paper. The PTR paradigm consists of three components:
@@
    AlphaEvolve makes an LLM call using the tuple of inputs in the prompt to the LLM, prompting it to output the description of $\mathcal{C}_{t+1}$, with the goal of increasing the score [^12] Technically, AlphaEvolve does not feed in all the prior functions as the prompt to the LLM. Instead, it uses a genetic algorithm to select a subset of them while maintaining high scores, and diversity [54, Figure 1].. We emphasize the following subtle point: AlphaEvolve does not directly search for combinatorial structures, but instead for a code snippet that would output a combinatorial structure with a high score. Figure 6 shows an initial code-snippet to be evolved by AlphaEvolve.
@@
    Our contributions: Our main contributions to the PTR paradigm are the following: a) using AlphaEvolve itself to accelerate the verification for $\textup{MAX-}{k}\textup{-CUT}$ (sometimes up to $10,000\times$), and b) an approach towards using AlphaEvolve for solving optimization problems that involve mixed-integer programming (MIP) style constraints. For the $k=3$ case, the AlphaEvolve optimized verification code constructed a verifier that offloaded a $O(3^{m})$ time computation to a highly optimized tensor contraction operation in numpy, and only performing slow python-based computation for $O(3^{m/3})$ time. We discussed each of these contributions in detail in Sections 3.1 and 3.2.
@@ Appendix B Proofs of Average Case Theorems
For all $d\in\mathbb{N}$ and $\varepsilon>0$, with probability $1-o_{n}(1)$ over $G\sim\mathcal{G}(n,d)$, we have $\lambda^{*}(G)\leq 2\sqrt{d-1}+\varepsilon$.
@@
The $d$-ary tree of depth $L$, $T_{d,L}$ is a rooted tree, where the root has $d$ children, and all non-root vertices at distance <L<L from the root have d1d-1 children.
@@
We will deduce some linear constraints on $\mu_{G,S}$, in terms of its probability distribution function.
@@
An interesting point is that the $d=4,L=3$ case has $11880$ variables and is technically within reach of current LP solvers, but resulted in the same bound on $\sigma^{\textup{MC}}_{4}$ as the $d=4,L=2$ case.
@@
In this section, we give the details of the NP-hardness of approximation proofs for MAX-k-CUT for $k\in\{3,4\}$. For ease of exposition, we begin by reintroducing all predicates needed for the proof. All predicates will be over the alphabet $\mathbb{Z}_{k}$.
  • $\texttt{P}^{\neq}_{2}$ denotes a binary predicate defined by $\texttt{P}^{\neq}_{2}(x,y)=\mathbf{1}\{x\neq y\}$.
  • For $i\in\{0,\ldots,k-1\}$, $\texttt{P}^{\equiv i}_{3,k}$ denotes a $3$-ary predicate defined by $\texttt{P}^{\equiv i}_{3,k}(x,y,z)=\mathbf{1}\{x+y+z\equiv i\pmod{k}\}$.
Recall that MAX-k-CUT is $\textup{CSP}(\texttt{P}^{\neq}_{2})$, and $3\textup{LIN}({k})$ is $\textup{CSP}\left(\texttt{P}^{\equiv 0}_{3,k},\ldots,\texttt{P}^{\equiv(k-1)}_{3,k}\right)$. We now formally define gadgets, specifically in the context of the reduction from $3\textup{LIN}({k})$ to $\textup{MAX-}{k}\textup{-CUT}$. We will need $k$ separate gadgets. In particular, for each defining predicate $\texttt{P}^{\equiv i}_{3,k}$ of $3\textup{LIN}({k})$, we will need an $i$-gadget, which we define below.
@@
Let $k\geq 2$.
  • $i$-gadget: For every $i\in\mathbb{Z}_{k}$, we introduce a gadget $\texttt{I}^{\equiv i}_{3,k}$ that reduces the predicate $\texttt{P}^{\equiv i}_{3,k}$ to an instance of $\textup{MAX-}{k}\textup{-CUT}$. Formally, $\texttt{I}^{\equiv i}_{3,k}$ is an instance of $\textup{MAX-}{k}\textup{-CUT}$ over $3+k+n_{\textup{aux}}$ variables, where $n_{\textup{aux}}\in\mathbb{N}$.
  • Primary, global, and auxiliary variables: We partition the set of variables $3+k+n_{\textup{aux}}]$ into three primary variables $\{1,2,3\}$, $k$ global variables $\{4,5,\ldots,k+3\}$, and $n_{\textup{aux}}$ auxiliary variables $\{k+4,\ldots,3+k+n_{\textup{aux}}\}$. We will always write an assignment to the variables as a tuple $(\mathbf{x},\mathbf{y},\mathbf{z})\in\mathbb{Z}_{k}^{3+k+n_{\textup{aux}}}$ where $\mathbf{x}\in\mathbb{Z}_{k}^{3}$, $\mathbf{y}\in\mathbb{Z}_{k}^{k}$, and $\mathbf{z}\in\mathbb{Z}_{k}^{n_{\textup{aux}}}$.
The variables $\mathbf{y}$ will be restricted in a way to break the symmetries of $\textup{MAX-}{k}\textup{-CUT}$. Concretely, let $Y$ be the set of strings $\mathbf{y}$ in $\mathbb{Z}_{k}^{k}$ that are lexicographically not larger than any string $\mathbf{y}^{\sigma}=(\sigma(y_{1}),\sigma(y_{2}),\ldots,\sigma(y_{k}))$ that is obtained by permuting the alphabet $\mathbb{Z}_{k}$ by some permutation $\sigma:\mathbb{Z}_{k}\to\mathbb{Z}_{k}$. In particular, note that the vector $(0,1,2,\ldots,k-1)$ is in $Y$. We will only ever assign $\mathbf{y}$ to values in $Y$.
@@
We will set $C_{i}=\left(\texttt{P}^{\equiv i}_{3,k}\right)^{-1}(1)$ and $S_{i}=\left(\texttt{P}^{\equiv i}_{3,k}\right)^{-1}(0)$ to be the set of satisfying and unsatisfying assignments for the source predicate $\texttt{P}^{\equiv i}_{3,k}$ respectively.
We will associate four real-valued parameters to the gadget $\texttt{I}^{\equiv i}_{3,k}=\sum_{j\in[m]}\zeta_{j}$ as follows.
  • (Completeness analysis) $c(\texttt{I}^{\equiv i}_{3,k})$ is defined as
    $c(\texttt{I}^{\equiv i}_{3,k})=\min_{\mathbf{x}\in C_{i}}\max_{\mathbf{z}\in\mathbb{Z}_{k}^{n_{\textup{aux}}}}\texttt{I}^{\equiv i}_{3,k}(\mathbf{x},\mathbf{y},\mathbf{z})$
    for $\mathbf{y}=(0,1,\ldots,k-1)$. For $\mathbf{x}\in C_{i}$, we say that $\arg\max_{\mathbf{z}\in\mathbb{Z}_{k}^{n_{\textup{aux}}}}\texttt{I}^{\equiv i}_{3,k}(\mathbf{x},\mathbf{y},\mathbf{z})$ is the witness for $\mathbf{x}$.
  • (Soundness analysis) $s_{1}(\texttt{I}^{\equiv i}_{3,k})$ is defined as
    $s_{1}(\texttt{I}^{\equiv i}_{3,k})=\max_{\mathbf{x}\in C_{i},\mathbf{z}\in\mathbb{Z}_{k}^{n_{\textup{aux}}},\mathbf{y}\in Y}\texttt{I}^{\equiv i}_{3,k}(\mathbf{x},\mathbf{y},\mathbf{z}).$
    Also, $s_{2}(\texttt{I}^{\equiv i}_{3,k})$ is defined as
    $s_{2}(\texttt{I}^{\equiv i}_{3,k})=\max_{\mathbf{x}\in S_{i},\mathbf{z}\in\mathbb{Z}_{k}^{n_{\textup{aux}}},\mathbf{y}\in Y}\texttt{I}^{\equiv i}_{3,k}(\mathbf{x},\mathbf{y},\mathbf{z}).$
  • (Number of clauses) $\texttt{count}(\texttt{I}^{\equiv i}_{3,k})$ is defined to be the number of clauses present in $\texttt{I}^{\equiv i}_{3,k}$.
@@
Let $k\geq 2$. Given an instance $\texttt{I}=\sum_{j\in[m]}\zeta_{j}$ of $3\textup{LIN}({k})$ with $m/k$ clauses from each of the predicates $\texttt{P}^{\equiv 0}_{3,k},\texttt{P}^{\equiv 1}_{3,k},\ldots,\texttt{P}^{\equiv(k-1)}_{3,k}$, it is NP-hard to distinguish between the following cases for any $\varepsilon>0$:
@@
    Completeness case. $\max_{\mathbf{x}}\texttt{I}(\mathbf{x})\geq(1-\varepsilon)\cdot m$, or
@@
    Soundness case. $\max_{\mathbf{x}}\texttt{I}(\mathbf{x})\leq(1/k+\varepsilon)\cdot m$.
@@
Let $k\geq 2$. Let $\texttt{I}^{\equiv i}_{3,k}$ be an $i$-gadget that reduces the predicate $\texttt{P}^{\equiv i}_{3,k}$ to an instance of $\textup{MAX-}{k}\textup{-CUT}$ for each $i\in\mathbb{Z}_{k}$. For any $\varepsilon>0$, given an instance J of $\textup{MAX-}{k}\textup{-CUT}$ with $M$ constraints, it is NP-hard to distinguish between the following two cases for any $\varepsilon>0$:
@@
    Completeness case: $\max_{\mathbf{x}}\texttt{J}(\mathbf{x})\geq(a-\varepsilon)\cdot M$, or
@@
    Soundness case: $\max_{\mathbf{x}}\texttt{J}(\mathbf{x})\leq(b+\varepsilon)\cdot M$,
@@
$a=\frac{\sum_{i\in\mathbb{Z}_{k}}c(\texttt{I}^{\equiv i}_{3,k})}{\sum_{i\in\mathbb{Z}_{k}}\texttt{count}(\texttt{I}^{\equiv i}_{3,k})},\quad b=\frac{s_{1}(\texttt{I}^{\equiv r_{0}}_{3,k})+\sum_{i\in\mathbb{Z}_{k}\setminus\{0\}}s_{2}(\texttt{I}^{\equiv r_{i}}_{3,k})}{\sum_{i\in\mathbb{Z}_{k}}\texttt{count}(\texttt{I}^{\equiv i}_{3,k})}.$
Here $r_{0},r_{1},\ldots r_{k-1}$ is a permutation of $\{0,1,\ldots,k-1\}$ such that $s_{1}(\texttt{I}^{\equiv r_{i}}_{3,k})-s_{2}(\texttt{I}^{\equiv r_{i}}_{3,k})$ is maximized by $r_{0}$.
@@
We will perform a reduction starting from the NP-hard problem in Theorem C.1 to $\textup{MAX-}{k}\textup{-CUT}$. Let $\texttt{I}=\sum_{j\in[m]}\zeta_{j}$ be an $n$-variable instance of $3\textup{LIN}({k})$, and suppose that it has $m/k$ clauses using $\texttt{P}^{\equiv i}_{3,k}$ for all $i\in\mathbb{Z}_{k}$. We will show how to reduce I to an instance J of $\textup{MAX-}{k}\textup{-CUT}$.
Let $n_{\textup{aux}}$ be the number of auxiliary variables in all gadgets $\{\texttt{I}^{\equiv i}_{3,k}:i\in\mathbb{Z}_{k}\}$; we can assume they all have the same number of auxiliary variables by introducing new dummy variables to some of them. $J$ will be defined as an instance on $n+k+n_{\textup{aux}}\cdot m$ variables, which we will denote as $\mathbf{x}\in\mathbb{Z}_{k}^{n}$, $\mathbf{y}\in\mathbb{Z}_{k}^{k}$, and $\mathbf{z}\in\mathbb{Z}_{k}^{n_{\textup{aux}}\cdot m}$. For $j\in[m]$ we will denote the $j^{th}$ block of size $n_{\textup{aux}}$ in $z$ as $z^{j}$. We will refer to $\mathbf{x},\mathbf{y},\mathbf{z}$ as the primary, global, and auxiliary variables respectively.
Let $\zeta_{j}=\texttt{P}^{\equiv i}_{3,k}(x_{j_{1}},x_{j_{2}},x_{j_{3}})$ be a clause corresponding to the predicate $\texttt{P}^{\equiv i}_{3,k}$ for some $i$. We will add a copy of $\texttt{I}^{\equiv i}_{3,k}$ on the variables $(\mathbf{x}^{\prime},\mathbf{y},\mathbf{z}^{j})$, where $\mathbf{x}^{\prime}=(x_{j_{1}},x_{j_{2}},x_{j_{3}})$. Note that the number of clauses in J is $M=(m/k)\cdot\sum_{i\in\mathbb{Z}_{k}}\texttt{count}(\texttt{I}^{\equiv i}_{3,k})$.
@@
    Completeness case: Assuming there is an $\mathbf{x}\in\mathbb{Z}_{k}^{n}$ such that $\texttt{I}(\mathbf{x})\geq(1-\varepsilon)\cdot m$, we show how to construct an assignment $(\mathbf{x},\mathbf{y},\mathbf{z})\in\mathbb{Z}_{k}^{n+k+n_{\textup{aux}}\cdot m}$ such that $\texttt{J}(\mathbf{x},\mathbf{y},\mathbf{z})\geq a\cdot M$. We will set $\mathbf{y}=(0,1,2,\ldots,k-1)$, and for each clause $\zeta_{j}=\texttt{P}^{\equiv i}_{3,k}(x_{j_{1}},x_{j_{2}},x_{j_{3}})$, $\mathbf{z}^{j}$ will be set as the witness
    $\mathbf{z}^{j}=\arg\max_{\mathbf{z}\in\mathbb{Z}_{k}^{n_{\textup{aux}}}}\texttt{I}^{\equiv i}_{3,k}(\mathbf{x}^{\prime},\mathbf{y},\mathbf{z})$, where $\mathbf{x}^{\prime}=(x_{j_{1}},x_{j_{2}},x_{j_{3}})$.
@@
    $\sum_{i\in\mathbb{Z}_{k}}c(\texttt{I}^{\equiv i}_{3,k})\cdot m/k-O(\varepsilon\cdot m)=a\cdot M-O(\varepsilon)\cdot m\geq(a-\varepsilon^{\prime})\cdot M$